OpenStudy (anonymous):

solve by stating the equation. after receiving a discount of 12.5% on its bulk order of printer paper,johns office supply pays $3200.what was the price of the order before the discount? round to nearest dollar.

OpenStudy (anonymous):

Hello agn. Sooo, after the discount the ORiginal price is reduced as follows:

OpenStudy (anonymous):

(1-0.125)* Original = 3200

OpenStudy (anonymous):

sorry (1 -0.0125)* Or =3200

OpenStudy (anonymous):

Now lets look for this O (original price)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

teach me please

OpenStudy (anonymous):

0.9875*O=3200

OpenStudy (anonymous):

O = 3200/0.9875

OpenStudy (anonymous):

?

OpenStudy (anonymous):

3240

OpenStudy (anonymous):

3240

OpenStudy (anonymous):

Gooood Giirls :)

OpenStudy (anonymous):

yay!!!!!!!

OpenStudy (anonymous):

So U found it urself

OpenStudy (anonymous):

yes wow i cant believe that lol...thanks to u taking the time to show me

OpenStudy (anonymous):

Fan (you have to somehow become "fan" of me in this mechanism I dont know how

OpenStudy (anonymous):

wow me neither this is my first time on here

OpenStudy (anonymous):

i did it

OpenStudy (anonymous):

put your mouse on my user name it brings up a box to the right will be become a fan